OpenStudy (anonymous):

Solve the problem. Be sure to show all of your work. The volume of a sphere with radius r is given by the formula V=4/3 πr3 Find the volume of a sphere with radius 3 meters. Use 3.14 for the value of π.

OpenStudy (anonymous):

if r = 3 meters (radius is 3 meters) and the volume of a sphere is V=4/3 πr^3 then all you need to do is plug 3 in for the value of r in that equation it would be: V=4/3 πr^3 V=4/3 π(3)^3 V = (27*4/3)π V = 36π